### v1

# Investigation of the effects of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere using spatial dynamic mode decomposition

Abstract
Viscous drag results from the fluid at a surface having zero relative velocity, a phenomenon known as the no-slip condition. Superhydrophobic surfaces, when submerged in water, trap a layer of air in their surface texture, partially replacing the liquid-solid interface with a liquid-gas interface. This air layer, called the plastron, results in partial slip at the surface, thereby reducing the viscous drag. In turbulent flows, large fluctuations in pressure and velocity can deplete or completely remove the plastron from the surface. This makes evaluating the effects of superhydrophobic surface treatments on flow dynamics particularly challenging. This study examines the impact of a sustained plastron on flow dynamics in the wake of a sphere, achieved by supplying air at low pressure through pores in the sphere’s surface. Instantaneous planar velocities in the wakes of spheres, both with and without superhydrophobic surface treatment, are measured within a plane passing through the sphere’s centre. Dynamic mode decomposition is applied to the velocity fluctuations in the wake to evaluate how superhydrophobic surface treatments affect flow dynamics in the near wake. It is shown that the addition of the pores has a relatively small effect on the dynamics in the wake of the sphere, while the dynamics in the wake are radically changed by the addition of superhydrophobic surface treatment when the plastron is sustained.
1 Introduction
1.1 Superhydrophobic Surfaces
The lotus leaf exhibits a unique ability to repel water due to a combination of surface roughness and chemical properties, enabling water droplets to roll off without wetting the surface [1, 2]. This phenomenon, termed superhydrophobicity, has garnered significant interest for its potential applications, particularly in drag reduction and anti-fouling technologies [3, 4]. Water droplets on a superhydrophobic surface form with a large contact angle, which is the internal angle between the surface and the edge of the droplet [5, 6]. When sliding down an inclined surface, the droplets exhibit low contact angle hysteresis, which is the difference between the angle of the droplet at its leading and trailing edges [7, 3]. The texture of superhydrophobic surfaces consists of multi-scale roughness, which amplifies the effect of chemical hydrophobicity [8]. Due to this combination of properties, the droplets are easily perturbed [9, 10, 11, 12].
These properties are achieved by trapping a layer of air, known as the plastron, between the surface and the droplet in what is known as the Cassie-Baxter state [13]. When submerged in water, the air trapped at the surface leads to a partial air-water interface and thus a partial slip over the surface [3], which reduces the skin friction at the surface and decreases the viscous drag of flow over the surface. A major factor in the varying results of experiments with superhydrophobic surfaces in turbulent flows is the propensity for the plastron to be depleted or completely deteriorated by the large velocity and pressure fluctuations present in turbulent flows. Even in cases where the plastron remains attached to the surface, these fluctuations have a significant effect on its distribution over the surface [14]. Once the plastron is depleted, the surface roughness begins to protrude into the flow, and the drag increases. This can occur across the entire surface as the plastron deteriorates, or in local regions as the plastron is deformed by the forces exerted on it by the flow.
As a plastron-covered superhydrophobic surface decreases viscous drag in water flows, it is of particular interest in reducing fuel consumption in marine vessels, such as ships and submarines, and for reducing transport costs in water supply pipes. While the drag reduction of superhydrophobic surfaces in laminar flows is directly related to the amount of slip produced, which is determined by the fraction of the surface aerated by the plastron [3], drag reduction in turbulent flows is more complicated to determine. While drag reduction has been achieved in turbulent flows, such as boundary layers [15] and Taylor-Couette flows [16], investigations of superhydrophobic drag reduction have also resulted in negligible changes [17], or even significant increases [18] in drag. The depletion of the plastron in turbulent boundary layers at high Reynolds number in the study by Aljallis et al. [19] resulted in an increase in drag, while the reduction of drag on a boat in open water observed by Xu et al. [20] relied on the plastron being sustained. The study of turbulent boundary layers on superhydrophobic surfaces by Gose et al. [21], over a Reynolds number range of 10,000 to 30,000, resulted in drag reduction of 90% to drag increases of 90%. This depended on the Reynolds number and surface properties, with the product of the surface-to-projected area ratio of the wetted region and the proportion of the surface area that is aerated being the best predictor of drag reduction.
1.2 Flow Structures in the Wake of A Sphere
As the application of superhydrophobic surfaces for drag reduction is of particular interest for maritime vessels, such as ships and submarines, the effects of superhydrophobic surface treatment on the flow over bluff bodies are an important consideration to assess the efficacy of these surfaces. Spheres serve as an idealised model for axisymmetric bluff bodies, exhibiting reverse flow, flow separation, and periodic and oscillatory vortex shedding in their wakes [22, 23, 24]. Quantitative and qualitative measurements in the wake of the sphere have identified numerous structures and characterised their development over various Reynolds numbers. The wake loses axisymmetry around $Re_{D}\approx 20$, with a vortex ring forming at $Re_{D}\approx 25$ [22, 25]. The vortex ring begins to oscillate at $Re_{D}\approx 130$, and elongates as the Reynolds number increases. Single-frequency vortex shedding begins at $Re_{D}\approx 300$ [26, 27, 24], and the generation of a lateral force [28, 29] results in the loss of planar symmetry by $Re_{D}\approx 500$. Small-scale vortical structures are present within the main vortex ring from $Re_{D}\approx 1,000$ as the vortex rings are shed from the sphere and break down as they move downstream [29, 30].
The drag of a sphere in motion in a fluid at rest can be determined using settling sphere experiments, in which the sphere is released from rest and accelerates under gravity until it reaches its terminal velocity $w_{\infty}$. The corresponding Reynolds number and drag coefficient [31] are given by
$Re_{\infty}=\frac{w_{\infty}d}{\nu}$ (1)
and
$C_{D,\infty}=\frac{4dg(\zeta-1)}{3w_{\infty}^{2}}$ (2)
respectively, where $d$ is the sphere diameter, $\nu$ is the kinematic viscosity of water, $g$ is the gravitational acceleration, and $\zeta$ is the sphere-to-fluid density ratio. In order to measure unsteady drag prior to the sphere reaching terminal velocity, the acceleration of the sphere is needed. Where the displacement is experimentally measured, the measurement noise in the acceleration, which is amplified by differentiation with a small time-step, can be significantly reduced using physics-informed neural networks [32].
The drag effects of superhydrophobic surfaces on spheres have been measured with this method using spheres with and without superhydrophobic surface treatment to determine the change in the terminal drag coefficient [33, 34, 35]. The terminal Reynolds number can be changed by using spheres of different diameters and densities [31, 34], which allows the Reynolds number effects on drag to be investigated. Observations of the plastron formed over the sphere during these experiments show significant changes to its distribution over the surface [14]. The plastron is deformed by the flow around the sphere such that it is more concentrated towards the rear of the sphere. This results in a feedback loop between the flow in the wake of the sphere and its plastron, with the plastron distribution around the back of the sphere changing with the instantaneous pressure fluctuations, which affect the flow around the sphere. Thus, the drag effects of superhydrophobic surface treatment on bluff body flows are not only a result of the change in conditions at the surface, but also the dynamic effects that the plastron has on the flow.
1.3 The Current Work
The current work investigates the effect of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere, with the plastron actively maintained by supplying air through pores in the surface of the sphere. Three spheres were manufactured: (i) a smooth untreated sphere to serve as a reference, (ii) an untreated sphere with a porous shell to determine the change in dynamics from the addition of the pores, and (iii) a porous sphere with superhydrophobic surface treatment to determine the change in the flow dynamics when the plastron is sustained.
The instantaneous velocities in a plane through the near wake of each sphere were measured using particle image velocimetry (PIV), from which the instantaneous velocity fluctuations were determined. Spatial dynamic mode decomposition (DMD) was performed in the streamwise direction on the velocity fluctuations in the near wake of each sphere. The resulting DMD modes were sorted from slowest to fastest decay rate. The slowest decaying DMD modes for each sphere were compared to determine the effect of the superhydrophobic surface treatment on the dynamics of the near wake when the plastron is maintained.
The results show that the addition of pores to the sphere has relatively little effect on the slowest-decaying DMD modes present in the near wake, although there are changes to the decay rates and wavelengths of these modes. In contrast, the addition of superhydrophobic surface treatment radically changes the DMD modes in the wake due to the distribution of the plastron over the sphere. While the three slowest-decaying DMD modes in the wakes of the untreated spheres exhibit planar symmetry, only the slowest-decaying DMD mode of the sphere with superhydrophobic surface treatment is planar symmetric. The effects of the increase in three-dimensionality in the flow on the plastron are likely to be highly dependent on the flow geometry and Reynolds number, and will affect the efficacy of the superhydrophobic surface treatment at reducing drag. Varying geometries and flow conditions are also likely to have a substantial effect on how the addition of superhydrophobic surface treatments affects the dynamics of the flow, which may enhance or mitigate the drag reduction effects of the partial slip introduced by the superhydrophobic surface.
2 Methodology
2.1 Experimental Facility
2.1.1 Vertical Water Tunnel
The flow in the near wake of the spheres was measured in a vertical water tunnel facility with a 250 mm $\times$ 250 mm cross-section and a 1.5 m working section [36, 37, 38], as shown in figure 1a. The settling chamber above the working section includes a 16:1 contraction and honeycomb screens to reduce the turbulence in the working section. The 0.5 m high test section is located in the middle of the working section, with a cut-out section on one wall to provide access. An acrylic panel is used to seal the tunnel, and is flush with the internal side of the wall to minimise the disturbance to the flow. Water is pumped from the plenum chamber below the working section up to the settling chamber through a return pipe, such that the flow through the working section is downwards. The introduction of seeding to the flow and the filtration of the tunnel are performed using an auxiliary circuit connected to the return pipe, which is isolated from the tunnel during experiments.
A crossbeam structure of thin aerofoils is mounted above the test section to support the sting to which the sphere is attached. The sting has a diameter of 9.3 mm until the last 100 mm, over which it tapers linearly to 4.65 mm and ends in a threaded rod for mounting the sphere. The sting and the supporting aerofoil structure include an internal channel connected to a compressed air supply, through which air was supplied to the superhydrophobic (SHS) sphere to maintain the plastron. A close-up view of the crossbeam structure, sting and sphere is shown in figure 2.
Figure 1: (a) Vertical water tunnel facility [38] and (b) experimental field of view.
Figure 2: Aerofoil crossbeam and sting used for mounting the sphere [38].
2.1.2 Sphere Design
The spheres were 3-D printed with a resolution of 22 µm $\times$ 22 µm in the horizontal plane and 35 µm in the vertical direction. The spheres were printed with a diameter of $D=40$ mm and a 2.5 mm-thick shell. An internal support structure was included to support the sphere during printing and the experiments, with slits to allow air to flow from the sting to the surface of the sphere. An external boss was included on the top of the sphere to smooth the transition from the sting to the sphere.
In order to allow air to reach the surface of the sphere, pores with a diameter of 1 mm were placed on the sphere’s surface according to the angular position given by
$\begin{bmatrix}\theta\\ \phi\end{bmatrix}_{i}=\begin{bmatrix}\frac{2\pi i}{\Phi_{GR}}\\ \\ \cos^{-1}\left(\frac{1-2(i+\frac{1}{2})}{N_{P}}\right)\end{bmatrix},$ (3)
where $\theta$ and $\phi$ are the azimuthal and polar angles, respectively, $N_{P}$ is the total number of pores over the sphere, and $\Phi_{GR}=\frac{1+\sqrt{5}}{2}$ is the golden ratio [38]. The number of points was chosen to achieve a porosity of 20%. The pores that would interfere with the external boss were removed, as were those positioned further than 50 from the top of the sphere. The latter condition was chosen to prevent escaping bubbles from reflecting laser light into the camera. A sphere with no pores or surface treatment was printed to serve as a reference (REF), and a porous (PRS) sphere was printed and left without surface treatment to determine the effect of the pores on the flow in the absence of superhydrophobic surface treatment. A second porous sphere was treated with Rust-Oleum NeverWetTM SHS coating to determine the effect of the sustained plastron on the flow. The spheres are shown in figure 3.
Figure 3: Cut-away view of 3D-printed sphere models for the (a) smooth reference sphere and (c) porous spheres with and without superhydrophobic surface treatment [38].
2.2 Experimental Method
The freestream velocity of the tunnel was set to $U_{\infty}=200$ mm/s for all spheres, resulting in a Reynolds number of $Re_{D}=7,780$, with a corresponding background turbulence of 2% [38]. The mean convective time-scale arising from the sphere diameter and freestream velocity is $t_{c}=0.2$[38]. The flow was seeded with glass spheres of 11µm diameter, with the single-exposed particle images illuminated by a pair of 400 mJ Spectra Physics Nd:YAG lasers firing 7 ns pulses and recorded using a PCO Panda 26 DS camera. The camera consists of a 5120 $\times$ 5120 array of 2.5 µm $\times$ 2.5 µm pixels, which was used to capture a 190 mm $\times$ 190 mm field of view beginning directly below the sphere and aligned to its centre, as shown in figure 1b, with a resulting spatial resolution of 37 µm/px. Pairs of single-exposed images were captured with a separation of $\Delta t=1$ ms, and image pairs were acquired at a rate of 1 Hz, resulting in a separation between instantaneous velocity fields of $5t_{c}$. The acquisition of $N_{T}>10,000$ images for each sphere was synchronised with the laser pulses using a BeagleBone Black (BBB) pulse generator [39].
The plastron formed on the SHS sphere was maintained by supplying air at low pressure, with the pores cleared of any settled seeding particles periodically [38]. The flow rate required to maintain the plastron without causing air to jet from the pores was 0.18$\pm$0.006 litres per minute. No roughness was added to the sphere, and as a result the plastron was thinner than those in settling sphere experiments with added roughness [14], as shown in figure 4. This reduced the risk of reflections into the camera in the field of view.
Figure 4: Image of the plastron formed on the superhydrophobic sphere, taken with a non-PIV camera [38].
Measurement of the instantaneous streamwise and transverse velocities, $u(x,y,t)$ and $v(x,y,t)$, respectively, in the wake of each sphere was performed using two-component-two-dimensional (2C-2D) multigrid/multipass cross-correlation digital (MCCD)-PIV [38, 40]. In order to balance the intensity of the single-exposed images, the mean background images of the first and second exposures were subtracted from the corresponding images before processing. The final interrogation window was 64 px $\times$ 32 px, with a grid spacing of 32 px $\times$ 16 px in $x$ and $y$, respectively. The instantaneous velocity fluctuations were determined for each vector field by subtracting the mean velocity field from the instantaneous velocity fields. The key experimental parameters are summarised in table 1.
Table 1: Table of experimental parameters [38].
Parameter Symbol Physical Units Relative Units
Freestream velocity $U_{\infty}$ 200 mm/s --
Sphere diameter $D$ 40 mm --
Reynolds number $Re_{D}$ 7,780 --
Time between laser pulses $\Delta t$ 1 ms $0.005t_{c}$
Time between velocity fields $\Delta T$ 1 s $5t_{c}$
Spatial resolution $SR$ 37 µm/px 9.25×$\times 10^{-4}D$
[…]
Number of 2C-2D velocity fields $N_{T}$ $>$10,000
The uncertainty of the 2C-2D-MCCDPIV algorithm used is 0.06 px [40], which corresponds to a single-sample uncertainty of 2.22 mm/s in the instantaneous velocity $u$, due to random error. The mean of 10,000 instantaneous velocity vectors, $\overline{u}$, therefore has an uncertainty of 0.0222 mm/s. As the instantaneous velocity fluctuations are given by
$u^{\prime}=u-\overline{u},$ (4)
their uncertainty $\sigma_{u^{\prime}}$ is given by
$\sigma_{u^{\prime}}^{2}=\sigma_{u}^{2}+\sigma_{\overline{u}}^{2},$ (5)
and is equal to 1.1% of the freestream velocity.
2.3 Dynamic Mode Decomposition
While turbulence often appears random, many turbulent flows contain coherent structures. These structures are not always evident in the instantaneous measurements of the flow; however, they play a vital role in processes such as heat transfer and mixing, as well as affecting the dynamic behaviour of aerodynamic and hydrodynamic forces, such as lift and drag. The identification of coherent structures and their effect on the flow can be used to create more parsimonious models of complex fluid dynamics problems by isolating those structures that have the greatest effect on the quantity of interest. Thus, identifying coherent structures in turbulent flows is pertinent to many practical applications of fluid dynamics.
Modal decomposition is commonly used to determine the time information in patterns in turbulent fluctuations, which may indicate coherent structures and their time-scale in the flow. The basis of modal decomposition is the singular value decomposition (SVD), which is also the basis of proper orthogonal decomposition (POD[41, 42]. The SVD of the $m\times n$ matrix $\boldsymbol{X}$, where $\boldsymbol{X}_{m,n}$ represents a scalar field at time $t_{m}$ and spatial coordinate $\boldsymbol{x}_{n}$, is given by
$\boldsymbol{X}=\boldsymbol{U}\Sigma\boldsymbol{V}^{T},$ (6)
where $\boldsymbol{U}$ and $\boldsymbol{V}^{T}$ contain the left and right eigenvectors of $\boldsymbol{X}$, respectively, and $\Sigma$ contains the corresponding singular values. For applications to turbulent flows, $\boldsymbol{X}$ is populated with the turbulent fluctuations, with the square of each singular value representing the turbulent kinetic energy of the corresponding row of $\boldsymbol{V}^{T}$. Each row of $\boldsymbol{V}^{T}$ is an orthonormal spatial mode of $\boldsymbol{X}$, with temporal coefficients given by the corresponding column of $\boldsymbol{U}$. The modes are sorted from largest to smallest $\Sigma$ and are thus ranked from the modes that contribute the most turbulent kinetic energy to $\boldsymbol{X}$ to those that contribute the least.
DMD is used to extract the dynamic characteristics of the matrix $\boldsymbol{A}$, such that
$\boldsymbol{X}_{t+\Delta t}=\boldsymbol{A}\boldsymbol{X}_{t},$ (7)
where $\boldsymbol{X}_{t}$ is a vector field at time $t$ and $\boldsymbol{X}_{t+\Delta t}$ is the vector field at time $t+\Delta t$, and the matrix $\boldsymbol{A}$ describes the dynamical processes relating the vector field to the subsequent vector field, which is an approximation of the Koopman operator [43]. Direct calculation and manipulation of $\boldsymbol{A}$ is computationally cumbersome for large data sets. Thus, exact DMD uses the reduced SVD of $\boldsymbol{X}_{t}$ to define the matrix
$\boldsymbol{\widetilde{A}}=\boldsymbol{U}^{T}_{r}\boldsymbol{X}_{t+\Delta t}\boldsymbol{V}_{r}\Sigma^{-1}_{r},$ (8)
where $r$ denotes the reduced-order SVD [44]. This is used to compute the DMD modes and their eigenvalues without requiring direct computation or manipulation of $\boldsymbol{A}$.
In order for DMD to be meaningful, the $\Delta t$ between vector fields must be adequately small to capture the dynamics of the flow from which the vector fields are taken, which limits its application to time-resolved [45] or pairwise time-resolved data [46]. When the vector fields are not separated by sufficiently small $\Delta t$ for DMD to be performed in time, the matrices $\boldsymbol{X}_{t}$ and $\boldsymbol{X}_{t+\Delta t}$ can be replaced with $\boldsymbol{X}_{x}$ and $\boldsymbol{X}_{x+\Delta x}$, respectively, where $\Delta x$ is the separation in the streamwise direction [43]. Thus, the dynamical processes being represented by the DMD are
$\boldsymbol{X}_{x+\Delta x}=\boldsymbol{A}\boldsymbol{X}_{x},$ (9)
where $\boldsymbol{A}$ represents the dynamical processes between the streamwise positions $x$ and $x+\Delta x$.
The matrix $\boldsymbol{X}_{x}$ was defined using the instantaneous velocity fluctuations at all streamwise locations except for the furthest downstream, where each row represents an instant in time and each column represents a position in space. In order for both velocity components to be included simultaneously, the matrix $\boldsymbol{X}_{x}$ was defined as
$\boldsymbol{X}_{x}=\begin{bmatrix}u^{\prime}\big{(}{(x,y)}_{1},t_{1}\big{)}&\cdots&u^{\prime}\big{(}{(x,y)}_{N},t_{1}\big{)}&v^{\prime}\big{(}{(x,y)}_{1},t_{1}\big{)}&\cdots&v^{\prime}\big{(}{(x,y)}_{N},t_{1}\big{)}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\\ u^{\prime}\big{(}{(x,y)}_{1},t_{N_{T}}\big{)}&\cdots&u^{\prime}\big{(}{(x,y)}_{N},t_{N_{T}}\big{)}&v^{\prime}\big{(}{(x,y)}_{1},t_{N_{T}}\big{)}&\cdots&v^{\prime}\big{(}{(x,y)}_{N},t_{N_{T}}\big{)}\\ \end{bmatrix},$ (10)
where $t_{m}$ is the time of the $m^{th}$ velocity field, ${(x,y)}_{n}$ is the $n^{th}$ spatial coordinate, and $N$ denotes the total number of spatial coordinates excluding those at the furthest downstream location. The matrix $\boldsymbol{X}_{x+\Delta x}$ was thus defined by replacing $x$ with $x+\Delta x$ in equation (10).
The SVD of $\boldsymbol{X}_{x}$ was computed using equation (6) and the reduced-rank SVD was defined as
${(\boldsymbol{X}_{x})}_{r}=\boldsymbol{U}_{r}\Sigma_{r}\boldsymbol{V}^{T}_{r},$ (11)
where $r$ is the rank of the reduction, and the rank-reduction ratio is given by
$RRR=\frac{\Sigma_{r}^{2}}{\Sigma_{1}^{2}}.$ (12)
A rank-reduction ratio of $RRR=0.25$ was chosen for all spheres to limit the DMD to the most dominant POD modes. This resulted in ranks of ten, nine, and seven for the REF, PRS, and SHS spheres, respectively. The cumulative and relative turbulent kinetic energy of the included POD modes are shown in figure 5 [38].
Figure 5: (a) Cumulative turbulent kinetic energy of the leading POD modes and (b) the ratio of the energy of the leading POD modes to the first POD mode for each sphere [38]. Included modes are shown by solid shapes and excluded modes are shown by hollow shapes.
The reduced-rank SVD of $\boldsymbol{X}_{x}$ and the full matrix $\boldsymbol{X}_{x+\Delta x}$ were then used to determine the matrix $\boldsymbol{\widetilde{A}}$ using equation (8). The eigendecomposition of $\boldsymbol{\widetilde{A}}$
$\boldsymbol{\widetilde{A}}\boldsymbol{\Phi}=\widetilde{\mu}\boldsymbol{\Phi},$ (13)
yields eigenvalues $\widetilde{\mu}$ and eigenvectors $\boldsymbol{\Phi}$, from which the DMD modes of $\boldsymbol{\widetilde{A}}$ are given by
$\psi_{k}(x,y)=\widetilde{\mu}_{k}{\boldsymbol{\Phi}}^{-1}_{k}\Sigma_{r}\boldsymbol{V}_{r}^{T}.$ (14)
where $k$ is an integer from 1 to $r$. The eigenvalues of $\boldsymbol{A}$, $\mu$, are related to those of $\boldsymbol{\widetilde{A}}$ by
$\mu_{k}=\frac{\log(\widetilde{\mu}_{k})}{\Delta x}.$ (15)
The growth/decay rate of the $k^{th}$ DMD mode in the streamwise direction, $r^{x}_{k}$, is given by the real part of $\mu_{k}$
$r^{x}_{k}=\Re(\mu_{k})=\frac{\log(|\widetilde{\mu}_{k}|)}{\Delta x},$ (16)
and the wavenumber is given by the imaginary part
$\Im(\mu_{k})=\frac{\theta_{\widetilde{\mu}_{k}}}{\Delta x},$ (17)
where $|\widetilde{\mu}_{k}|$ and $\theta_{\widetilde{\mu}_{k}}$ are the magnitude and angle of $\widetilde{\mu}_{k}$ in the complex plane, respectively. The wavenumber defined in equation (17) is in radians per metre and is divided by $2\pi$ to be converted to m-1, therefore the wavelength of the $k^{th}$ DMD mode in the streamwise direction, $\lambda^{x}_{k}$ is
$\lambda_{k}^{x}=\frac{2\pi}{\Im(\mu_{k})}=\frac{2\pi\Delta x}{\theta_{\widetilde{\mu}_{k}}}.$ (18)
The eigenvalues of $\boldsymbol{A}$ and $\boldsymbol{\widetilde{A}}$ are shown in figures 6a and 6b, respectively. The decay rates and wavelengths of the three slowest-decaying DMD modes are given in table 2, and the corresponding DMD modes in the wake of each sphere are shown in figures 78, and 9, respectively. In order to provide an additional visualisation of the structures in the DMD modes, the out-of-plane vorticity
$\omega^{\prime}(x,y)=\frac{\partial v^{\prime}(x,y)}{\partial x}-\frac{\partial u^{\prime}(x,y)}{\partial y},$ (19)
of the real and imaginary parts of each DMD mode was calculated. The resulting vorticity fields in the wake of each sphere are shown in figures 1011, and 12, for the first, second, and third DMD modes of each sphere, respectively.
3 Results
Figure 6: Eigenvalues of (a) $\widetilde{\boldsymbol{A}}$ and (b) $\boldsymbol{A}$ from spatial DMD. The black dashed line in (a) is the unit circle.
Table 2: Decay rates $r^{x}_{k}$, per diameter, and wavelengths $\lambda^{x}_{k}$, in diameters, of the first three DMD modes in the wakes of the REF, PRS and SHS spheres.
DMD Mode, $k$ 1 2 3
REF $r^{x}_{k}$0.07 -0.08 -0.10
$\lambda^{x}_{k}$ 4.48 2.27 1.95
PRS $r^{x}_{k}$0.06 -0.07 -0.08
$\lambda^{x}_{k}$ 4.52 2.22 2.40
SHS $r^{x}_{k}$0.06 -0.11 0.13
$\lambda^{x}_{k}$ 3.08 9.70 22.67
All DMD modes resulting from the spatial dynamic mode decomposition have negative real parts, as shown in figure 6b, meaning that they all decay in the streamwise direction. This is expected, as turbulent flows are typically stable and do not feature convective instabilities. The first two DMD modes of the REF and PRS spheres have very similar decay rates and wavenumbers, while there is a significant difference in the decay rates and wavenumbers of their third DMD modes. This suggests that the two DMD modes with the slowest decay rates are unaffected by the addition of the pores, while the pores have some effect on the dynamics that decay more quickly in the streamwise direction. The decay rates and wavenumbers of the first three DMD modes in the wake of the SHS sphere are significantly different from those of the untreated spheres, suggesting that the addition of superhydrophobic surface treatment has a significant effect on the dynamics in the near wake of the sphere. This is consistent with the leading DMD modes of the proper orthogonal decomposition, compared in [38], which are very similar for the untreated spheres but radically different for the SHS sphere.
Figure 7: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the first DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The first DMD modes are qualitatively similar between the REF and PRS spheres, as shown in figures 7a and 7b, respectively. The streamwise components of the real part of the first DMD modes for each sphere exhibit antisymmetry across $y=0$, with regions of alternating sign in the streamwise direction, as shown in figures 7a(i) and 7b(i), respectively. The corresponding transverse components are symmetric across $y=0$, also with regions of alternating sign in the streamwise direction, as shown in figures 7a(ii) and 8b(ii), respectively. This behaviour is reflected in the imaginary parts of the streamwise components, as shown in figure 7a(iii) and 7b(iii), respectively, and the transverse components, shown in figure 7a(iv) and 7b(iv), respectively, which show the same structures propagating downstream. In addition to the qualitative similarity, the slowest decaying DMD modes of the REF and PRS spheres have comparable wavelengths of $4.48D$ and $4.52D$, respectively. The decay rate of the first DMD mode of the PRS sphere decays slightly more slowly than that of the REF sphere, at $-0.06D^{-1}$ and $-0.07D^{-1}$, respectively. Although qualitatively similar in structure to the first DMD modes of the REF and PRS spheres and decaying at a similar rate of $-0.06D^{-1}$, the first DMD mode of the SHS sphere, shown in figure 7c, has a significantly shorter wavelength of $3.08D$, which is reflected in the reduced length of the structures in the SHS mode compared to the REF and PRS modes.
Figure 8: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the second DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The second DMD modes of the REF and PRS spheres, shown in figures 8a and 8b, respectively, are qualitatively similar to each other and have comparable decay rates, $-0.08D^{-1}$ and $-0.07D^{-1}$, and wavelengths, $2.27D^{-1}$ and $2.22D^{-1}$, respectively. Additionally, the structure of these modes is similar to that of the first mode but with the length of the structures in the streamwise direction reduced, which is reflected in the wavelengths being decreased by a factor of two compared to those of the first modes of each sphere. As with the first DMD modes of these spheres, the imaginary part of the second DMD modes, the streamwise and transverse components of which are shown in figures 8a(iii) and 8b(iii), and figures 8a(iv) and 8b(iv), respectively, represent the propagation of the real parts, the streamwise and transverse components of which are shown in figures 8a(i) and 8b(i), and figures 8a(ii) and 8b(ii), respectively, in the streamwise direction. The second DMD mode in the SHS case, shown in figure 8c, is substantially different from those of the untreated spheres. This is particularly evident when comparing the streamwise component of the second DMD mode, where the positive region directly behind the sphere, shown in figure 8c(i), becomes substantially larger and moves off centre as the negative region at $y/D=0.5$ is replaced with asymmetric negative regions at $y/D=\pm 0.5$, shown in figure 8c(iii).
Figure 9: Real (i) streamwise and (ii) transverse, and imaginary (iii) streamwise and (iv) transverse components of the third DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The third DMD modes of the REF and PRS spheres are also qualitatively similar, although the third PRS mode, shown in figure 9b, is significantly less symmetric than the third REF mode, shown in figure 9a. The asymmetry is likely due to the spherical distribution of pores on the PRS sphere, as the surface of the sphere on either side of the plane measured may not have exactly the same distribution of pores, resulting in an uneven boundary layer across the sphere. In addition to the introduction of asymmetry, the PRS mode has a lower decay rate of $-0.08D^{-1}$ compared to $-0.10D^{-1}$ for the REF sphere, and a longer wavelength of $2.40D$ compared to $1.95D$. This is reflected in the size of the structures, which are larger in the PRS case than in the REF case. As in the second DMD mode, the third DMD mode of the SHS sphere, shown in figure 9c, exhibits large structures that have substantial transverse variation in place of the streamwise propagation seen in the modes of the untreated spheres.
Figure 10: (i) Real and (ii) imaginary out-of-plane vorticity of the first DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The periodicity in the streamwise direction of the first DMD modes in the wake of each sphere is seen clearly in their out-of-plane vorticity, shown in figure 10. In addition to the reduced wavelength of the structures in the wake of the SHS sphere seen in the velocity components, the vorticity contour, shown in figure 10c, shows a planar asymmetry where the untreated spheres are planar symmetric, as shown in figures 10a and 10b. The growth of the wake can also be seen in these structures, whose transverse extent increases in the streamwise direction.
Figure 11: (i) Real and (ii) imaginary out-of-plane vorticity of the second DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The out-of-plane vorticity contours of the second DMD mode in the wakes of the REF and PRS spheres, shown in figures 11a and 11b, respectively, show similar structures to the first DMD modes but with reduced wavelength. This is consistent with the velocity components of these modes, which were also qualitatively similar. The vortical structures in the second DMD mode of the SHS sphere, shown in figure 11c, are substantially different from the untreated spheres. However, the positive vorticity directly behind the sphere is bounded by thinner negative regions with the sign of these regions changing after $x/D=3$, which is similar to the structures seen in the first and second DMD modes of the untreated spheres. Unlike the REF and PRS cases, the imaginary part of the second SHS mode, shown in figure 11c(ii), does not resemble a streamwise translation of its real part, shown in figure 11c(i). Although the large positive region in the imaginary part is bounded by negative regions, the pattern is heavily biased in the transverse direction, and the streamwise extent of the positive region is significantly longer than in the real part. An additional negative region is also present directly behind the sphere, within the mean recirculation region.
Figure 12: (i) Real and (ii) imaginary out-of-plane vorticity of the third DMD mode in the wake of the (a) REF, (b) PRS and (c) SHS spheres. The mean recirculation region is outlined in black.
The out-of-plane vorticity of the third DMD mode in the REF sphere, shown in figure 12a, is planar antisymmetric with structures directly behind the sphere bounded by structures of opposing sign. As with the velocity components, shown in figure 9a, the intensity of the out-of-plane vorticity decays in the streamwise direction. The out-of-plane vorticity of the third DMD mode in the PRS sphere, shown in figure 12b, shows the same structures elongated in the streamwise direction, with the planar asymmetry seen in the velocity, shown in figure 9b, components reflected in the out-of-plane vorticity. The out-of-plane vorticity of the third DMD mode of the SHS case, shown in figure 12c, shows no streamwise periodicity, as was the case with the velocity components. The imaginary part of the out-of-plane vorticity, shown in figure 12c(ii), appears to represent a rotation of the structures in the real part, shown in figure 12c(i), around the streamwise axis, instead of a propagation in the streamwise direction, as seen in the third DMD modes of the untreated spheres.
4 Discussion
The shapes of the first and second DMD modes of the REF and PRS spheres, and the first mode of the SHS sphere, are similar and consistent with the flapping of the wake. The second DMD modes of the REF and PRS cases are harmonics of the corresponding first DMD modes, as the wavelengths of the second DMD modes are approximately half those of the first DMD modes. The third DMD modes of the REF and PRS cases represent a pulsating motion, as alternating positive and negative streamwise velocity fluctuations propagate downstream. The first dynamic mode of the SHS sphere represents a flapping motion of the wake, as is the case for the untreated spheres. However, the wavelength and decay rate of this fluid motion are significantly altered by the superhydrophobic surface treatment.
The second and third DMD modes in the SHS case are substantially different from those of the REF and PRS cases. There is a more significant change in decay rate between the first DMD mode and the second and third DMD modes in the SHS case than in the case of the untreated spheres. Additionally, the wavelengths of the second and third DMD modes in the SHS case are significantly longer than the measurement domain, which is $4.7D$ in the streamwise direction. This suggests that these DMD modes may not be periodic in the streamwise direction, but they extend far enough downstream of the sphere that this cannot be verified with the current measurement domain. This is reflected in the DMD mode shapes, which do not exhibit the streamwise patterns observed in the other DMD modes. Despite the substantial distortion, the basic vortical structure of the real part of the second DMD mode of the SHS case is similar to that of the first SHS mode and the first and second DMD modes in the REF and PRS cases. The third DMD mode of the SHS sphere shows significantly more substantial three-dimensional effects than the third DMD modes of the REF and PRS spheres, with its real and imaginary parts representing a rotation around the streamwise axis rather than a propagation in the streamwise direction, as seen in the DMD modes of the untreated spheres.
While the leading DMD modes in the wake of the untreated spheres all exhibit a degree of planar symmetry, only the first DMD mode in the SHS case shows any degree of planar symmetry. The second and third DMD modes in the wake of the SHS sphere show more significant three-dimensionality than those of the untreated spheres. The reduction in planar symmetry in the DMD modes is most likely due to the changing shape of the plastron as it is subjected to pressure fluctuations across the sphere, as observed in the wakes of spheres with superhydrophobic surface treatment settling under gravity [14]. As the plastron is deformed by the fluid flow, the changing shape of the plastron affects the flow in turn, resulting in a feedback loop like that proposed in [47], which makes the flow significantly more complex dynamically than that in the wake of the untreated spheres. This added complexity, which would vary significantly with different geometries and flow conditions, is consistent with the highly variable drag effects seen in previous investigations of superhydrophobic surface treatments [19, 21].
5 Concluding Remarks
The streamwise and transverse velocities in a plane in the near wake of spheres with and without superhydrophobic surface treatment were measured using 2C-2D MCCD-PIV, with a large number of instantaneous velocity fields. The mean planar velocities were then determined and used to calculate the instantaneous velocity fluctuations, which were decomposed into DMD modes by applying dynamic mode decomposition in the streamwise direction. The resulting DMD modes were sorted by decay rate, from slowest to fastest, and the first three DMD modes for each sphere were compared.
The first two DMD modes in the REF and PRS cases were found to be qualitatively and quantitatively similar, with matching DMD mode shapes, decay rates, and wavelengths. The third DMD modes in these cases were qualitatively similar, albeit with the PRS case being somewhat asymmetric relative to the REF case. This asymmetry is likely due to the spherical distribution of the pores on the sphere resulting in slightly different flow conditions on either side of the sphere in the measured plane. While the first mode of the SHS sphere was qualitatively similar to those of the REF and PRS cases, the wavelength was significantly shorter. The second and third DMD modes of this case were radically different from the untreated spheres, suggesting a significant influence from the plastron behaviour on the dynamics in the near wake, greater three-dimensionality in the wake of the SHS sphere than in the REF and PRS cases.
The current work shows that the addition of superhydrophobic surface treatment has a significant effect on the dynamics of the flow in the near wake of a sphere, which is consistent with the variation in drag effects reported in previous studies. Thus, the drag effects of superhydrophobic surface treatment in turbulent flow are not just a function of the partial slip generated at the surface when the plastron is maintained, but also of the effect that the plastron has on the flow as it is deformed by the velocity and pressure fluctuations.
Acknowledgements
Shaun Davey gratefully acknowledges the support of the Australian Commonwealth Government through a Research Training Program (RTP) Scholarship.
This research was undertaken with the support of resources provided via a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government.